OpenStudy (anonymous):

What is the inequality? 4|2w+3|-7<9 I know the answer is 3.5

OpenStudy (anonymous):

To solve the inequality, you have to isolate the expression within the absolute value symbol first

OpenStudy (anonymous):

But what does that mean?

OpenStudy (anonymous):

You have to get the \[\left|2w+3? \right|\] by itself

OpenStudy (anonymous):

\[\left| 2w+3 \right|\]

OpenStudy (anonymous):

I've been doing that but it comes out as w<4.5 instead of w<3.5

OpenStudy (anonymous):

Could you show your work?

OpenStudy (anonymous):

Sure, And I actually meant it's suppose to be w<-3.5

OpenStudy (anonymous):

4|2w+3|-7<9 |2w+3|-3<9 -Add 3 to both sides- 2w+3<12 -Subtract 3- 2w<9 w<4.5 And I try to solve it different every time but it still doesn't come out right.

OpenStudy (anonymous):

And I did this way too. 4|2w+3|-7<9 -9<4(2w+3)-7<9 -9<8w+12-7<9 Add 5 -14<8w<4 Divide 8 -7/4<w<1\2

OpenStudy (anonymous):

I think what you're trying to do is combine the 4 and the -7, which you can't do since the 4 is the coefficient of the absolute value

OpenStudy (anonymous):

So are you saying I need to do distributive property? Because I did that in the second one.

OpenStudy (anonymous):

This is how the problem should be worked out. \[4\left| 2w+3 \right|-7<9\] \[4\left| 2w+3 \right|<16\] we added the 7 to both sides \[\left| 2w+3 \right|<4\] divided both sides by 4 Now we set up 2 inequalities \[2w+3<4\] \[2w+3>-4\]

OpenStudy (anonymous):

and is the answer really 3.5<w<.5 or did you mean to put -3.5<w<.5

OpenStudy (anonymous):

Yeah, I added earlier that I mean -3.5. Thank you so much! That makes a lot more sense now!

OpenStudy (anonymous):

You're welcome.